A bowl contains candies of the same size in three flavors: orange, strawberry, and pineapple. If the probability of randomly pulling out an orange candy is 1 over 9, and the probability of randomly pulling out a strawberry candy is 3 over 9, what is the probability of randomly pulling out a pineapple candy?
To find this add 3 and 1 together. This gives you 4. Then, you subtract 4 from 9. The answer is 5. Therefore, the probability to pull out a pineapple candy would be 5/9.
The probability of randomly pulling out a pineapple candy is:
\(\dfrac{5}{9}\)
There are three flavors in a ice cream namely:
orange, strawberry, and pineapple.
We know that for any three events in a single experiment the sum of the probability of three events is equal to 1.
i.e. if A, B and C are three events and P denotes the probability of an event then:
P(A)+P(B)+P(C)=1
Here let A= flavor is orange
B= flavor is strawberry
and C= flavor is pineapple.
and we have:
\(P(A)=\dfrac{1}{9},\ P(B)=\dfrac{3}{9}\)
Hence, we have:
\(\dfrac{1}{9}+\dfrac{3}{9}+P(C)=1\\\\\\i.e.\\\\\\\dfrac{4}{9}+P(C)=1\\\\\\i.e.\\\\\\P(C)=1-\dfrac{4}{9}\\\\\\i.e.\\\\\\P(C)=\dfrac{9-4}{9}\\\\i.e.\\\\P(C)=\dfrac{5}{9}\)